### v1

@@ 1 Introduction
The volume mesh-based methods, which discretize an analysis domain into finite volume elements, are widely used in structural optimization. Specifically, the topology optimization method has the advantage of optimizing structures with a high degree of freedom. After the pioneering work by Bendsøe and Kikuchi (1988), topology optimization has been applied to various applications in the structural mechanics field, for example, aeroplane wings (Aage et al, 2017), automotive components (Kim et al, 2020), and drone frames made of composites (Zhou et al, 2022b). The basic concept of topology optimization is to replace the original structural optimization problem by a material distribution problem in a fixed design domain such that arbitrary shapes can be represented. Usually, the topology optimization results in nonuniform thickness, but in the case of shell structures, they must have a certain plate thickness due to the raw material and manufacturing constraints. Therefore, topology optimization methods with geometric constraints have been studied extensively. Guest (2009) introduced the maximum length scale control in topology optimization. Allaire et al (2016) proposed a framework to constrain the local thickness via a level set-based method. Zhou et al (2022a) proposed a method for optimizing 3D thin-walled structures through two PDE-based filtering operations and an aggregation constraint. Such methods can be used to optimize shell-like structures while maintaining a near-uniform thickness. A coating approach (Clausen et al, 2015), which represents thin structures in the transition region between void and solid (infill), has also been used for shell structure optimization. The coating approach has mainly been applied to the shells with infills such as porous infill (Wang et al, 2018) or graded lattice (Wu et al, 2017). Clausen et al (2017) discussed assigning void to the interior, representing pure shell structures.
@@
Topology optimization on the shells has also been studied. It is often combined with shape optimization, both parametric methods (Ansola et al, 2002; Hassani et al, 2013; Kang and Youn, 2016; Jiang et al, 2023) and nonparametric methods (Shimoda et al, 2021; Ho-Nguyen-Tan and Kim, 2022). Simultaneous optimization of shape and topology has gained attention in recent years as it can increase design freedom by considering material distribution on the surface. However, both parametric and nonparametric shape optimization in their methods still suffer from the aforementioned problems for significant design changes of the surface itself.
@@ 2.2 Filtering and projection scheme
where $R$ is the filtering radius, $\mathbf{n}$ is the normal vector on design domain boundaries, and $\psi$ is a scalar function. $\psi$ is the design variable in the context of the optimization problem.
@@ 2.3 Design update of shell structure
where $c$ is a constant that should be set to $1/\|\nabla\phi\|$.
@@ 2.4 Shell analysis
In this study, computation of shell analyses and their sensitivities is conducted by using MITC6 triangular 6-node shell finite elements implemented in COMSOL Multiphysics (Domínguez Alvarado and Díaz Díaz, 2018). The MITC stands for “Mixed Interpolation of Tensorial Components” which refers to a mathematical technique utilized for alleviating numerical problems such as shear and membrane locking. These issues often arise in conventional shell element formulations (Chapelle and Bathe, 2010). In the MITC formulation, the transverse shear strains are replaced with assumed shear strains which are derived from an alternative displacement interpolation and are then linked to the displacements at specific tying points. For detailed information on the exact interpolations and the tying points used in MITC6 shell elements, refer to Kim and Bathe (2009); Lee and Bathe (2010).
@@ 2.5 Optimization problem
where $W_{\mathrm{s}}$ is the elastic strain energy density and $G_{\mathrm{max}}$ is the prescribed maximum allowable volume.
@@ 3.2 Sensitivity embedding and filtering
For the design update of the shell structure, the sensitivity of the level set function given by Eq. 9 is required. Here, the sensitivity of the original shell offset and level set function are defined by surface and volume, respectively. Thus, the sensitivity field on the surface must be embedded into the volume domain. This can be accomplished by embedding the corresponding nodes from surface to volume because the conformal mesh, which is volumetric, contains the nodes of the surface mesh. Note that we allow that the absolute scale of the sensitivity field in volume differs depending on the discretization of the domain, e.g., the mesh size. Yet, the relative value of the sensitivity is consistent to the original shell offset sensitivity.
@@
where $\widetilde{\frac{\partial F}{\partial\phi}}$ is the smoothed sensitivity of level set function and $R_{\mathrm{sens}}$ is the filtering radius. Eq. (22) is introduced to avoid the change of structure on $\Gamma_{\mathrm{D}}$ through the optimization process.
@@ 4.1 Dome
We consider two types of geometric constraints in the dome example. In the first case, the optimization proceeds while maintaining the shape of the top edge, allowing changes to the bottom edge. The second case introduces stricter constraints that fix both the top and bottom edges, focusing on internal structural adjustments. The volume constraint is not applied in the dome problem. For both cases, the shell thickness is set to 0.01. The applied load is uniformly vertical and downward, with a magnitude of 1. The Young’s modulus is set to $10^{5}$. The maximum and minimum mesh size specified in remeshing are 0.04 and 0.025, respectively. The filtering radius $R$ and $R_{\mathrm{sens}}$ are set to 0.05. The constant $c$ is set to 15, which is an approximated value of $1/\|\nabla\phi\|$ in the initial solution. In the optimization process, the update rate of the steepest descent method is set to 100.
@@ 4.2 Plate
The problem setting for a plate example is shown in Fig. 8. The initial structure is a flat plate, which shows low stiffness due to low second moment of area. The design domain size is $1\times 1\times 0.5$, where the flat plate is placed at the middle in $z$-axis. The volume constraint and the geometric constraint are not applied to this problem. The applied load is uniformly vertical and downward, with a magnitude of 0.01. The shape of loaded edge can be changed through the optimization process. The Young’s modulus is set to $10^{5}$. The maximum and minimum mesh size specified in remeshing are 0.04 and 0.025, respectively. The filtering radius $R$ and $R_{\mathrm{sens}}$ are set to 0.05. The constant $c$ and the update rate of the steepest descent method are set to 15 and 100, respectively.
@@
The optimization results are shown in Fig. 9. The flat plate was gradually changed to a corrugated plate which uses the height of design domain as much as possible. As a result, the mean compliance was significantly decreased as shown in Fig. 10. Compared to the existing nonparametric methods, the proposed method is advantageous to maintain the mesh regularity through the optimization with large shape changes. The surface mesh at 300 iteration is not distorted in spite of the significant change from initial shape. These results indicate that the proposed method is suitable for the optimization with large shape changes.
@@ 4.3 Cantilever beam
The problem setting for a cantilever beam example is shown in Fig. 11. The initial structure is a cylindrical surface with the radius of 0.2. One end is fixed, and another end is loaded uniformly vertical and downward, with a magnitude of 1. The design domain size is $1\times 0.6\times 0.8$ where the cylindrical surface is placed at the center in $yz$ plane. The shape change of loaded edge is suppressed by a Dirichlet boundary condition. The Young’s modulus is set to $10^{5}$. The maximum and minimum mesh size specified in remeshing are 0.04 and 0.025, respectively. The filtering radius $R$ and $R_{\mathrm{sens}}$ are set to 0.05. The constant $c$ and the update rate of the steepest descent method are set to 15 and 100, respectively.
@@
In this example, the volume enclosed by the shell is constrained with $G_{\mathrm{max}}=0.13$. For the volume constraint, the initial parameter in barrier function $\gamma_{p}$ was set as follows:
@@
The optimization results of the cantilever beam problem are shown in Fig. 12. From 0 to 50 iteration, the shape extended in $z$-axis to increase the stiffness. The surfaces around the fixed edge were dented at the center of $z$-axis. From 50 to 100 iteration, the two adjacent surfaces fused together. As a result, a bifurcated structure was obtained from the single cylindrical surface. In other words, the topology of the structure was changed. The optimized structure at iteration 300 has arched ribs on the side, which are often used in sheet metal parts to increase stiffness. As shown in Fig. 13, the optimized structure shows lower mean compliance compared to the initial structure while satisfying the volume constraint of the enclosed domain.
@@ 5 Conclusion
For future work, incorporating constraints on the shell material volume, considering buckling effects, and integrating shell topology optimization on the surface are promising directions to improve the applicability of the method.

### v2

@@ 1 Introduction
The volume mesh-based methods, which discretize an analysis domain into finite volume elements, are widely used in structural optimization. Specifically, the topology optimization method has the advantage of optimizing structures with a high degree of freedom. After the pioneering work by Bendsøe and Kikuchi (1988), topology optimization has been applied to various applications in the structural mechanics field, for example, aeroplane wings (Aage et al, 2017), automotive components (Kim et al, 2020), and drone frames made of composites (Zhou et al, 2022b).
The basic concept of topology optimization is to replace the original structural optimization problem by a material distribution problem in a fixed design domain such that arbitrary shapes can be represented. Usually, the topology optimization results in nonuniform thickness, but in the case of shell structures, they must have a certain plate thickness due to the raw material and manufacturing constraints. Therefore, topology optimization methods with geometric constraints have been studied extensively.
Guest (2009) introduced the maximum length scale control in topology optimization. Allaire et al (2016) proposed a framework to constrain the local thickness via a level set-based method. Zhou et al (2022a) proposed a method for optimizing 3D thin-walled structures through two PDE-based filtering operations and an aggregation constraint. Such methods can be used to optimize shell-like structures while maintaining a near-uniform thickness. A coating approach (Clausen et al, 2015), which represents thin structures in the transition region between void and solid (infill), has been used for shell structure optimization. A level set-based method has also been applied for coated structures (Wang and Kang, 2018). In their work, the coating material is represented by a zero-neighborhood of the level set function. The coating approach has mainly been applied to the shells with infills such as porous infill (Wang et al, 2018) or graded lattice (Wu et al, 2017). Clausen et al (2017) discussed assigning void to the interior, representing pure shell structures.
@@
Topology optimization on the shells has also been studied. Park and Youn (2008) proposed a level set-based topology optimization method for a material distribution problem on shells. Huo et al (2022) applied a moving morphable method for topology optimization on complex surfaces using conformal mapping. It is often combined with shape optimization, both parametric methods (Ansola et al, 2002; Hassani et al, 2013; Kang and Youn, 2016; Jiang et al, 2023b) and nonparametric methods (Shimoda et al, 2021; Ho-Nguyen-Tan and Kim, 2022). Simultaneous optimization of shape and topology has gained attention in recent years as it can increase design freedom by considering material distribution on the surface. However, both parametric and nonparametric shape optimization in their methods still suffer from the aforementioned problems for significant design changes of the surface itself.
@@ 2.2 Filtering and projection scheme
where $R$ is the filtering radius, $\mathbf{n}$ is the normal vector on design domain boundaries, and $\psi$ is a scalar function. $\psi$ is the design variable in the context of the optimization problem. The filter radius $R$ controls the minimum length scale in the level set function. Accordingly, a larger $R$ leads to a smaller maximum curvature of shells. In the proposed method, a large enough $R$ is needed to ensure that the curvature of shells can be represented by the actual mesh size.
@@ 2.3 Design update of shell structure
where $c$ is a constant that should be set to representative value of $1/\|\nabla\phi\|$.
The approximation error of the sensitivity is composed of several factors, such as surface mesh extraction and sensitivity projection from the surface mesh to the fixed mesh, and affects the accuracy of the design update. One important factor for the accuracy of the design update is the gradient of the scalar field near the zero-value boundary.
In the proposed method, the spatial gradient $\nabla\phi$ at $\phi=0$ is implicitly controlled by a combination of filtering and projection (Eq. (2)–(8)); however, the uniformity of the spatial gradient is not guaranteed mainly because it depends on the geometry, e.g., curvature and neighboring distance of zero isosurface. If the actual spatial gradient $\nabla\phi$ were used in Eq. (9), the level set sensitivity derived from the shape sensitivity would be extremely large when $\|\nabla\phi\|$ is very small, and the design collapses by update. Assuming $c$ as constant ensures that the optimization remains stable even with large shape changes and topological changes where the linear expansion of the level set function may not hold, although the design update is not proportional to the shape sensitivity.
@@ 2.4 Shell analysis
In this study, computation of shell analyses and their sensitivities is conducted by using MITC6 triangular 6-node shell finite elements implemented in COMSOL Multiphysics (Domínguez Alvarado and Díaz Díaz, 2018). Note that shell elements other than MITC6 can be used in the proposed method because the level set-based structural representation only requires that the structure can be discretized into surface mesh. The MITC stands for “Mixed Interpolation of Tensorial Components” which refers to a mathematical technique utilized for alleviating numerical problems such as shear and membrane locking. These issues often arise in conventional shell element formulations (Chapelle and Bathe, 2010). In the MITC formulation, the transverse shear strains are replaced with assumed shear strains which are derived from an alternative displacement interpolation and are then linked to the displacements at specific tying points. For detailed information on the exact interpolations and the tying points used in MITC6 shell elements, refer to Kim and Bathe (2009); Lee and Bathe (2010).
@@ 2.5 Optimization problem
where $W_{\mathrm{s}}$ is the elastic strain energy density and $G_{\mathrm{max}}$ is the prescribed maximum allowable volume. Note that we do not apply a typical volume constraint that limits material usage. The proposed shell representation inherently avoids an obvious solution that is entirely filled with material, thus such a material usage constraint is not mandatory.
@@ 3.1 Mesh generation using Mmg
The mesh size should be enough fine for the filter radius $R$ to accurately represent the level set isosurface and ensure accurate shell analysis. We set the maximum allowed element size by -hmax option in Mmg to keep the mesh resolution.
@@ 3.2 Sensitivity embedding and filtering
For the design update of the shell structure, the sensitivity of the level set function given by Eq. (9) is required. Here, the sensitivity of the original shell offset and level set function are defined by surface and volume, respectively. Thus, the sensitivity field on the surface must be embedded into the volume domain. This can be accomplished by embedding the corresponding nodes from surface to volume because the conformal mesh, which is volumetric, contains the nodes of the surface mesh. Figure 3 shows the procedure of design sensitivity derivation. Note that we allow that the absolute scale of the sensitivity field in volume differs depending on the discretization of the domain, e.g., the mesh size. Yet, the relative value of the sensitivity is consistent to the original shell offset sensitivity.
@@
where $R_{\mathrm{sens}}$ is the filtering radius. $\widetilde{\frac{\partial F}{\partial\phi}}$ is the smoothed sensitivity and it is assumed to be the sensitivity of design variable $\frac{\partial F}{\partial\psi}$. Since this filter replaces the filter sensitivity calculation in standard topology optimization, which shows the smoothing effect with radius $R$ (Kawamoto et al, 2011), $R_{\mathrm{sens}}$ should be set comparable to $R$. Eq. (22) is introduced to avoid the change of structure on $\Gamma_{\mathrm{D}}$ through the optimization process.
@@ 4.1 Dome
We consider two types of geometric constraints in the dome example. In the first case, the optimization proceeds while maintaining the shape of the top edge, allowing changes to the bottom edge. The second case introduces stricter constraints that fix both the top and bottom edges, focusing on internal structural adjustments. The volume constraint is not applied in the dome problem. For both cases, the shell thickness is set to 0.01. The applied load is uniformly vertical and downward, with a magnitude of 1. The Young’s modulus is set to $10^{5}$. The mesh size of regular mesh is 0.05. The maximum and minimum mesh size specified in remeshing are 0.04 and 0.025, respectively. The filtering radius $R$ and $R_{\mathrm{sens}}$ are set to 0.05. The constant $c$ is set to 15, which is an approximated value of $1/\|\nabla\phi\|$ in the initial solution. In the optimization process, the update rate of the steepest descent method is set to 100.
@@ 4.2 Plate
The problem setting for a plate example is shown in Fig. 8. The initial structure is a flat plate, which shows low stiffness due to low second moment of area. The design domain size is $1\times 1\times 0.5$, where the flat plate is placed at the middle in $z$-axis. The volume constraint and the geometric constraint are not applied to this problem. The applied load is uniformly vertical and downward, with a small magnitude of 0.01 to ensure it complies with the assumption of infinitesimal deformations. The shape of loaded edge can be changed through the optimization process. The Young’s modulus is set to $10^{5}$ and the shell thickness is set to 0.01. The mesh size of regular mesh is 0.05. The maximum and minimum mesh size specified in remeshing are 0.04 and 0.025, respectively. The filtering radius $R$ and $R_{\mathrm{sens}}$ are set to 0.05. The constant $c$ and the update rate of the steepest descent method are set to 15 and 100, respectively.
@@
The optimization results are shown in Fig. 9. The flat plate was gradually changed to a corrugated plate which uses the height of design domain as much as possible. As a result, the mean compliance was significantly decreased as shown in Fig. 10.
The corrugated plate configuration demonstrates an increased stiffness against bending loads, consistent with the findings of previous studies on shell shape optimization (Shimoda and Liu, 2014). The proposed method, which can accommodate larger shape changes, has shown the potential to further improve bending stiffness. Compared to the existing nonparametric methods, the proposed method is advantageous to maintain the mesh regularity through the optimization with large shape changes. The surface mesh at 300 iteration is not distorted in spite of the significant change from initial shape. These results indicate that the proposed method is suitable for the optimization with large shape changes.
@@ 4.3 Cantilever beam
The problem setting for a cantilever beam example is shown in Fig. 11. The initial structure is a cylindrical surface with the radius of 0.2. One end is fixed, and another end is loaded uniformly vertical and downward, with a magnitude of 1. The design domain size is $1\times 0.6\times 0.8$ where the cylindrical surface is placed at the center in $yz$ plane. The shape change of loaded edge is suppressed by a Dirichlet boundary condition. The Young’s modulus is set to $10^{5}$ and the shell thickness is set to 0.01. The mesh size of regular mesh is 0.05. The maximum and minimum mesh size specified in remeshing are 0.04 and 0.025, respectively. The filtering radius $R$ and $R_{\mathrm{sens}}$ are set to 0.05. The constant $c$ and the update rate of the steepest descent method are set to 15 and 500, respectively.
@@
In this example, the volume enclosed by the shell is constrained with $G_{\mathrm{max}}=0.13$ to demonstrate the capability for handling geometric constraints. For the volume constraint, the initial parameter in barrier function $\gamma_{p}$ was set as follows:
@@
The optimization results of the cantilever beam problem are shown in Fig. 12. From 0 to 50 iteration, the shape extended in $z$-axis to increase the stiffness. The surfaces around the fixed edge were dented at the center of $z$-axis. From 50 to 100 iteration, the two adjacent surfaces fused together. As a result, a bifurcated structure was obtained from the single cylindrical surface. In other words, the topology of the structure was changed. The optimized structure at iteration 300 has arched ribs on the side, which are often used in sheet metal parts to increase stiffness. As shown in Fig. 13, the optimized structure shows lower mean compliance compared to the initial structure while satisfying the volume constraint of the enclosed domain. The shape change resulting in a thicker root section of the cantilever beam is similar to the vertical section shape in the T-shaped joint result of a previous study (Shimoda et al, 2021). However, due to the constraint of the enclosed volume in our case, the final shape differed substantially, including the topology.
@@ 5 Conclusion
For future work, incorporating constraints on the shell material volume, considering buckling effects are promising directions. In addition, addressing material distribution problems on shells and optimizing the layout of stiffeners (Jiang et al, 2022, 2023a), such as ribs, is crucial for industrial applications. Therefore, further advancements in structural representation methods and sophisticated mesh generation techniques are demanded to improve the applicability of the method.
Funding
No funding was received for conducting this study.
Statements and Declarations
[…]